1. A basketball has a radius of 12.5 inches. What would be the volume of the basketball?

Answer by chessace(471)

You need the formula for volume of a sphere:
V = 4/3 Pi r^3
Here V = 1.33 * 3.14 * 12.5^3
= 4.176 * 1953 = 8156 cubic inches.
Note that one could have carried more decimal places but they would not be justified, the precision of 12.5 and basketball charaterics limit such.